In an isosceles triangle, the base is 7 more than one-half times the legs. If perimeter is 22 centimeters. Find length of each leg.
Answer by lwsshak3(11628)

In an isosceles triangle, the base is 7 more than one-half times the legs. If perimeter is 22 centimeters. Find length of each leg.
**
let x=length of each leg
base+2 legs=perimeter
(1/2)x+7+2x=22
x/2+7+2x=22
5x/2=15
5x=30
x=6
length of each leg=6 cm